# Propagation of two-particle correlations across the chaotic phase for interacting bosons

### v1

@@ Abstract
We present a detailed analysis of the propagation of experimentally relevant two-particle correlations for one-dimensional interacting bosons, and give evidence that many-body chaos induces the emergence of an effective diffusive regime for the fully coherent correlation dynamics, characterized by an interaction dependent diffusion coefficient, which we estimate. This result supports very recent experimental observations, and paves the way towards an efficient description of the dynamical behaviour of non-integrable complex many-body systems. Furthermore, we show that the dynamical features within experimentally accessible time scales of a conveniently defined two-particle correlation transport distance provide a direct and unambiguous characterization of many-body quantum chaos.
Complex many-body dynamics ensue from an involved interplay between many-particle interference and interactions. The latter, in particular, are responsible for the appearance of chaotic regimes characterized by universal dictates [1, 2, 3], in which certain coarse-grained features of a many-body system become amenable to a description in terms of random matrix theory and statistical mechanics [4, 5, 6, 7, 8]. The specificity of the system in the chaotic phase, however, remains imprinted in certain observables [9, 11, 12], and the sensitivity to many-particle interference is in fact enhanced at the onset of such ‘universal’ regime [12, 13]. This rises a fundamental issue, how does the many-particle correlation dynamical response change upon the emergence of many-body quantum chaos?
@@
Here, we show compelling evidence that the existence of many-body quantum chaos for one-dimensional bosons triggers a fundamental change in the coherent dynamics of two-particle correlations, giving rise to an effective diffusive regime governed by an interaction dependent diffusion constant. This result is in significant agreement with very recent experimental observations for hard-core bosons [34]. We find that the time development of correlations can be encoded into a suitable two-particle correlation transport distance, whose behaviour —within experimentally reachable time scales— further permits an explicit identification of the chaotic phase in remarkable agreement with its spectral characterization [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 9, 10, 11, 12].
@@
An overview of the dynamical behaviour of $\ell(\tau)$ for $L=17$ is shown in Fig. 1(a). The CTD undergoes an initial growth in time and, due to the finiteness of the system, eventually reaches a $\gamma$-dependent saturation value which can be well identified for $\tau\in[100,200]$. The most interesting feature of the signals in this time interval is the striking absence of temporal fluctuations for some $\gamma$ values. The CTD saturation value, identified as the time average $\ell_{\text{sat}}\equiv\overline{\ell(\tau)}$ for $\tau\in[100,200]$, is shown in Fig. 1(b) as a function of $\gamma$. The propagation of correlations is strongly suppressed for $\gamma<0.1$, where $\ell_{\text{sat}}\ll 1$ and it exhibits a quadratic dependence on $\gamma$. As the relative tunneling strength further increases, $\ell_{\text{sat}}$ undergoes a pronounced surge, reaching a maximum followed by a subtle decay towards the non-interacting limit (see also Fig. S2 in the supplemental material). Our analysis shows that $\ell_{\text{sat}}$ exhibits a linear growth with $L$ for all values of $\gamma$ [59], $\ell_{\text{sat}}=aL+b$, albeit with markedly different slopes. In fact, around $\gamma=0.11$ a change in the dependence of $\ell_{\text{sat}}$ on $\gamma$ becomes visible for increasing $L$. This transition is distinctively witnessed by the fit parameters $a$ and $b$ [inset of Fig. 1(b)], suggesting that a non-analyticity develops in the vicinity of $\gamma=0.11$ in the thermodynamic limit.
@@
The question now is to which extent this dynamical regime correlates with the system’s spectrally chaotic phase. The latter is best identified from the eigenvector structure [9, 10, 11, 12], which is efficiently probed by the generalized fractal dimension $\tilde{D}_{1}=-\sum_{\ket{\boldsymbol{n}}}|\braket{\boldsymbol{n}}{\psi}|^{2}\log|\braket{\boldsymbol{n}}{\psi}|^{2}/\log\mathcal{D}$, for close-in-energy eigenstates |$\ket{\psi}$ expanded in the on-site Fock basis {|$\{\ket{\boldsymbol{n}}\}$. The top panel of Fig. 2 shows the energy resolved variance of $\tilde{D}_{1}$ for $L=11$ calculated over sets with 1% of the eigenstates, as a function of $\gamma$ and of the spectral percentage distance (sp) from the energy of the initial state |$\ket{\psi_{0}}$ taken as a reference value. In this stationary picture, as demonstrated in Refs. [9, 10, 11, 12], the chaotic phase is revealed by the region with suppressed $\operatorname{var}(\tilde{D}_{1})$, which correlates with the emergence of extended eigenstates ($\tilde{D}_{1}\to 1$) in Fock space in the thermodynamic limit [59]. We note that the onset of the chaotic domain so identified does not change with increasing $L$ [59]. To see how the initial state participates of the system eigenstates, the reference energy trajectory at sp $=0$ is flanked by two bands corresponding to ±$\pm\sigma\equiv\pm(\braket{H^{2}}_{\psi_{0}}-\braket{H}^{2}_{\psi_{0}})^{1/2}$, which gives the energy width of the local density of states in Fock space at the point |$\ket{\psi_{0}}$. The top panel of Fig. 2 thus shows how the spectrally chaotic phase develops as a function of $\gamma$ from the perspective of |$\ket{\psi_{0}}$, whose passage through the chaotic phase (accompanied by a marked broadening of the energy width) correlates unambiguously with the regime of suppressed temporal fluctuations of the CTD. Furthermore, $(\operatorname{var}_{\tau}\ell)/\ell_{\text{sat}}^{2}$ is minimized in the $\gamma$-range where the initial state participates maximally of eigenstates that exhibit the closest agreement with random matrix theory (RMT) (see contour line in the top panel of Fig. 2). These results also confirm that the dynamical emergence of ETH hinges on the existence of chaotic (extended in Fock space) eigenstates.
@@
We have presented a detailed analysis of the propagation of experimentally relevant two-particle correlations for one-dimensional interacting bosons. Our results show evidence that in the chaotic regime certain aspects of the otherwise complex many-particle fully coherent dynamics are amenable to a simplified approach using a diffusive model specified by an interaction dependent diffusion coefficient. This observation goes in hand with very recent experimental results for low-dimensional hard-core bosons [34], and paves the way towards an efficient description of the dynamical behaviour of non-integrable complex many-body systems. Furthermore, we showed that the time development of two-point density correlations can be conveniently encoded in a correlation transport distance, whose features within experimentally accessible time scales provide a direct and unambiguous characterization of many-body quantum chaos.

### v2

@@ Abstract
We analyze the propagation of experimentally relevant two-particle correlations for one-dimensional interacting bosons, and give evidence that many-body chaos induces the emergence of an effective diffusive regime for the fully coherent correlation dynamics, characterized by an interaction dependent diffusion coefficient, which we estimate. This result supports very recent experimental observations, and paves the way towards an efficient description of the dynamical behaviour of non-integrable complex many-body systems. Furthermore, we show that the dynamical features within experimentally accessible time scales of a conveniently defined two-particle correlation transport distance provide a direct and unambiguous characterization of many-body quantum chaos in perfect agreement with its spectral identification.
Complex many-body dynamics ensue from an involved interplay between many-particle interference and interactions. The latter, in particular, are responsible for the appearance of chaotic regimes characterized by universal dictates [1, 2, 3], in which certain coarse-grained features of a many-body system become amenable to a description in terms of random matrix theory and statistical mechanics [4, 5, 6, 7, 8]. The specificity of the system in the chaotic phase, however, remains imprinted in certain observables [9, 10, 11], and the sensitivity to many-particle interference is in fact enhanced at the onset of such ‘universal’ regime [12, 13]. This raises a fundamental issue, how does the many-particle correlation dynamical response change upon the emergence of many-body quantum chaos?
@@
Here, we show compelling evidence that the existence of many-body quantum chaos for one-dimensional bosons triggers a fundamental change in the coherent dynamics of two-particle correlations, giving rise to an effective diffusive regime governed by an interaction dependent diffusion constant. This result is in significant agreement with very recent experimental observations for hard-core bosons [34]. We find that the time development of correlations can be encoded into a suitable two-particle correlation transport distance, whose behaviour —within experimentally reachable time scales— further permits an explicit identification of the chaotic phase in remarkable agreement with its spectral characterization [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 9, 55, 10, 11]. Most interestingly, we show evidence that the equilibrium value of such correlation transport distance develops a non-analyticity (as the thermodynamic limit is approached) as a function of the interaction strength at the onset of the chaotic regime.
@@
An overview of the dynamical behaviour of $\ell(\tau)$ for $L=17$ is shown in Fig. 1(a). The CTD undergoes an initial growth in time and, due to the finiteness of the system, eventually reaches a $\gamma$-dependent saturation value which can be well identified for $\tau\in[100,200]$. The most interesting feature of the signals in this time interval is the striking absence of temporal fluctuations for some $\gamma$ values. The CTD saturation value, identified as the time average $\ell_{\text{sat}}\equiv\overline{\ell(\tau)}$ for $\tau\in[100,200]$, is shown in Fig. 1(b) as a function of $\gamma$. The propagation of correlations is strongly suppressed for $\gamma<0.1$, where $\ell_{\text{sat}}\ll 1$ and it exhibits a quadratic dependence on $\gamma$. As the relative tunneling strength further increases, $\ell_{\text{sat}}$ undergoes a pronounced surge, reaching a maximum followed by a subtle decay towards the non-interacting limit (see also Fig. S2 in the supplemental material). Our analysis shows that $\ell_{\text{sat}}$ exhibits a linear growth with $L$ for all values of $\gamma$ [59], $\ell_{\text{sat}}=aL+b$, albeit with markedly different slopes. In fact, around $\gamma=0.11$ a change in the dependence of $\ell_{\text{sat}}$ on $\gamma$ becomes visible for increasing $L$. This transition is distinctively witnessed by the fit parameters $a$ and $b$ [inset of Fig. 1(b)], indicating that a non-analyticity develops in the vicinity of $\gamma=0.11$ in the thermodynamic limit.
@@
The question now is to which extent this dynamical regime correlates with the system’s spectrally chaotic phase. The latter is best identified from the eigenvector structure [9, 55, 10, 11], which is efficiently probed by the generalized fractal dimension
$\tilde{D}_{1}=-\sum_{\ket{\boldsymbol{n}}}|\braket{\boldsymbol{n}}{\psi}|^{2}\log|\braket{\boldsymbol{n}}{\psi}|^{2}/\log\mathcal{D},$ (4)
for eigenstates |$\ket{\psi}$ expanded in the on-site Fock basis {|$\{\ket{\boldsymbol{n}}\}$. The top panel of Fig. 2 shows the energy resolved variance of $\tilde{D}_{1}$ for $L=11$ calculated over sets with 1% of the eigenstates, as a function of $\gamma$ and of the spectral percentage distance (sp) from the energy $E=0$ of the initial state |$\ket{\psi_{0}}$ taken as a reference value (i.e., the spectral percentage at energy $E$ corresponds to $sp=100\times[\mathcal{N}(E)-\mathcal{N}(0)]/\mathcal{D}$, with $\mathcal{N}(E)$ the integrated density of states).
In this stationary picture, as demonstrated in Refs. [9, 55, 10, 11], the chaotic phase is revealed by the region with suppressed $\operatorname{var}(\tilde{D}_{1})$, which correlates with the emergence of eigenstates that become extended ($\tilde{D}_{1}\to 1$) in Fock space in the thermodynamic limit [59]. We note that the onset of the chaotic domain so identified does not change with increasing $L$ [59]. To see how the initial state participates of the system eigenstates, the reference energy trajectory at sp $=0$ is flanked by two bands corresponding to ±$\pm\sigma\equiv\pm(\braket{H^{2}}_{\psi_{0}}-\braket{H}^{2}_{\psi_{0}})^{1/2}$, which gives the energy width of the local density of states in Fock space at the point |$\ket{\psi_{0}}$. The top panel of Fig. 2 thus shows how the spectrally chaotic phase develops as a function of $\gamma$ from the perspective of |$\ket{\psi_{0}}$, whose passage through the chaotic phase (accompanied by a marked broadening of the energy width) correlates unambiguously with the regime of suppressed temporal fluctuations of the CTD. Furthermore, $(\operatorname{var}_{\tau}\ell)/\ell_{\text{sat}}^{2}$ is minimized in the $\gamma$-range where the initial state participates maximally of eigenstates that exhibit the closest agreement with random matrix theory (RMT) (see contour line in the top panel of Fig. 2). These results also confirm that the dynamical emergence of ETH hinges on the existence of chaotic (extended in Fock space) eigenstates.
@@
We have studied the propagation of experimentally relevant two-particle correlations for one-dimensional interacting bosons. Our results show evidence that in the chaotic regime certain aspects of the otherwise complex many-particle fully coherent dynamics are amenable to a simplified approach using a diffusive model specified by an interaction dependent diffusion coefficient. This observation goes in hand with very recent experimental results for low-dimensional hard-core bosons [34], and paves the way towards an efficient description of the dynamical behaviour of non-integrable complex many-body systems. Furthermore, the time development of two-point density correlations can be conveniently encoded in a two-particle correlation transport distance (CTD), whose features within experimentally accessible time scales provide a direct and unambiguous characterization of many-body quantum chaos. Besides demonstrating analytically the ballistic growth of the CTD in the vicinity of the system’s integrable limits, we specifically show that (i) the onset of many-body chaos correlates with the emergence of a non-analyticity in the equilibrium value of the CTD when approaching the thermodynamic limit, and (ii) the regime of suppressed CTD temporal fluctuations expose the chaotic phase in a one-to-one correspondence with its spectral identification.